calculate the work done by the in twisting a steel wire of radius 1.22mm and the length 40cm through an angle 30 degree .the rigidity modulus of the material of wire r=4.55*10^10 N/m^2.

Grade:12

1 Answers

Nirmal Singh.
10 years ago
Work Done = energy = 1/2(c)*Theta^2
Twisting Constant c = pi eta r^4 /2 l
c = 3.14*4.55x10^10*(1.22x10^-3)^4 / 2*40x10^-2
c = 3.9563 = 4 (approx.)
So Work done = 1/2 *4*(30)^2 = 1800 Joule
Thanks & Best Regards,
Nirmal Singh
Askiitians faculty